OpenStudy (anonymous):

(____)^2 = (cscx - 1)(cscx + 1)

OpenStudy (anonymous):

fill in the blank.

OpenStudy (anonymous):

use the identity \[\csc ^2x-\cot ^2x=1\]

OpenStudy (anonymous):

I dont know how to.

OpenStudy (chris911):

istnt it (cot)^2

OpenStudy (anonymous):

\[\csc ^2x-1=\cot ^2x\]

OpenStudy (anonymous):

isn't what cot^2?

OpenStudy (anonymous):

\[\left( a+b \right)\left( a-b \right)=a^2-b^2\]

OpenStudy (anonymous):

a=csc x,b=1

OpenStudy (anonymous):

oh so the answer is cot^2?